A jar has two nickels, three quarters, and a half-dollar coin in it. Three coins are randomly drawn.

  1. (4) List the simple events in the sample space if order does not matter. You may use a code such as N = nickel 1, n = nickel 2, Q = quarter 1, q = quarter 2, p = quarter 3, and H = the half-dollar coin.
  2. (2) Show how to use a permutation or a combination formula to determine the number of options that could occur in the sample space if order does not matter.
  3. (2) Show how to use a permutation or a combination formula to determine the number of options that could occur in the sample space if order does matter.
  4. (1) What is the probability that a selection of three coins includes a half-dollar coin?
  5. (1) What is the probability that the selection of three coins does not include a half-dollar coin?
  6. (2) What is the probability that the selection of three coins is worth less than $1?
  7. (2) What is the probability that the selection of three coins is equal to $1?
  8. (2) What is the probability that the selection of three coins is greater than $1?
  9. (4) Draw or create a cumulative frequency table for the ways in which the sum of the coins can be shown. Hint: You want to show $0.35, $0.55, $0.60, $0.75, $1, etc. for the x column.

You may add or delete rows to the table below.

x

frequency

P(x)

Cumulative frequency

Cumulative P(X < x)

$0.35

$0.55

$0.60

$0.75

$1
Answer #1

All possible types of coins are: N, n, Q, q, p, H.

  • Samples of three can be chosen in: { (N,n,Q), (N,n,q), (N,n,p), (N,n,H), (N,Q,q), (N,Q,p), (N,Q,H), (N,q,p), (N,q,H), (N,p,H), (n,Q,q), (n,Q,p), (n,Q,H), (n,q,p), (n,q,H), (n,p,H), (Q,q,p), (Q,q,H), (Q,p,H), (q,p,H) }.
  • We consider each coins are distinct. There are 6 different kinds of coins. From the 6 coins, 3 coins can be chosen in \binom{6}{3} ways. Hence, combination is used to choose the coins. Permutation will not be used, as arrangements are not taken into consideration.

  • From the 6 coins, Half a dollar coin can be selected in 1 ways.
    The remaining 2 coins from the 5 coins can be selected in \binom{5}{2} ways.
    Hence, the required probability is =\frac{\binom{5}{2}*1}{\binom{6}{3}}=0.5

  • Without the half a dollar coin, there are 5 coins.
    Thus, from the remaining 5 coins, three coins can be selected in \binom{5}{3} ways.
    Hence, the required probability is =\frac{\binom{5}{3}*1}{\binom{6}{3}}=0.5 .

  • For the sum of the values of the coin to be less than $1, the combinations of one half cent and two quarter cents should not be taken into account.
    The half dollar coin can be chosen in 1 way and the two quarter dollar coin can be chosen in \binom{3}{2} ways.
    Hence, the required probability is =\frac{\binom{6}{3}-\binom{3}{2}*1}{\binom{6}{3}}=0.85 .

  • For the sum of the values of the coin to be equal to $1, only the combinations of one half cent and two quarter cents should be taken into account.
    The half dollar coin can be chosen in 1 way and the two quarter dollar coin can be chosen in \binom{3}{2} ways.
    Hence, the required probability is =\frac{\binom{3}{2}*1}{\binom{6}{3}}=0.15 .

  • It is observed that, P(sum of amount is less than $1)+P(sum of amount is equal to $1) =1.
    Hence,
    P(sum of amount is greater than $1)
    = 1 - P(sum of amount is less than or equal to $1)
    = 1 - (P(sum of amount is less than $1)+P(sum of amount is equal to $1))
    = 1 - 1
    = 0.

As per HOMEWORKLIB POLICY, we are to answer the first four subparts. However, I have tried answering as many questions as I could in the given time limit.

I hope this clarifies your doubt. If you're satisfied with the solution, hit the Like button. For further clarification, comment below. Thank You. :)